i need some basic rules for the following type of problem set:
(i^0) * (-i^0) = ???
does i^o * i^0 = 1 ???
-i * i = ???
does -i^0 = -1 ???
i^1 * -i^1 = ???
i^2 * -i^2 = ???
-i^1 * i^2 = ???
i^1 * -i^2 = ???
then the set:
i^-1 * i^-2 = ??? ; i^1 * 1^-2 = ??? ; etc.
any resources on this type of factoring would be extremely useful, since i haven't seen this expressed anywhere in the literature as of yet.

this type of factoring might be simple to roger penrose, but it's driving me nuts.

Answer by stanbon(75887)

Powers of i follow a cyclic pattern that you should memorize.
i^0=1
i^1=i
i^2=-1
i^3=-i
i^4=1
Then it starts over again because the last result was "1".
Following this pattern you should be able to work your problems.
Cheers,
Stan H.